OpenStudy (anonymous):

If 1 u.s dollar is .78 euros, how many euros will I have for 156 u.s dollars. Please show calculations. Thanks.

OpenStudy (anonymous):

111.808 Because 1 EUR = 1.39525 USD You can write the math

OpenStudy (anonymous):

1 us dollar/.78 euros = 156 us dollars/ E euros cross multiply: 1 * E = .78 * 156

OpenStudy (anonymous):

pi.i please shoe calculations

OpenStudy (anonymous):

cross multiply (1us dollar * E euros) = (.78 euros * 156 us dollars) then drop the labels since they are the same on both sides (1*E)=(.78*156) thus E = .78 *1 56 euros, (i don't have a calculator)

OpenStudy (anonymous):

what does that symbol * mean?

OpenStudy (anonymous):

pi.i awaiting reply

OpenStudy (anonymous):

sorry computer problem, * is to multiply